# Cross-Document Cross-Lingual NLI via RST-Enhanced Graph Fusion and Interpretability Prediction

### v2

@@ E.2 LLM Answer Case
Qwen3-8B Reasoning Output Okay, let’s tackle this NLI task. So, the premise has two articles in Spanish and Czech, and the hypothesis is in English. I need to determine if the hypothesis is entailed, contradicted, or neutral based on the premises. First, I’ll read through the Spanish article. It mentions that on Saturday, there was a fire in a gold mine in Condesuyos province, caused by a short circuit 100 meters from the entrance. The mine is called Yanaquihua, known as Esperanza I. They say 27 people died, two were rescued. The regional government of Arequipa mentioned the short circuit caused the fire, risking workers’ lives. Local media said 27 trapped workers died of asphyxia. The Ministry of Interior confirmed the accident on Saturday, and police were there for rescue. The fire spread quickly through wooden structures, preventing escape. The mine belongs to Yanaquihua S.A.C., a company with small miners. Now the Czech article: Officials said the accident in the small gold mine in southern Peru resulted in 27 deaths. It happened on Saturday morning at the Yanaquihua mine in Condesuyos province, Arequipa department. A short circuit caused a fire inside the tunnel. The regional government confirmed 27 victims. The prosecutor Giovanni Matos said the information is correct, and police are there to find the dead, check safety for rescue teams. Regional government and interior ministry mobilized police, medics, and ambulances. It’s the highest number of deaths in a single mining accident since 2000. The company didn’t respond to a comment request. Hypothesis: The Yanaquihua gold mine in Condesuyos, Peru, had a fatal incident on Saturday morning due to a short circuit causing a fire, leading to 27 workers’ deaths trapped inside the tunnel, with local authorities and rescue teams mobilized. Check each part. The Spanish article says the fire was caused by a short circuit, leading to 27 deaths. The Czech article confirms the same: short circuit causing fire, 27 deaths. Both mention Saturday morning (the Spanish says "sábado" which is Saturday, and the Czech says "v sobotu ráno" which is Saturday morning). The mine is Yanaquihua in Condesuyos, Peru. The death count is 27, and both mention authorities and rescue teams being mobilized. The hypothesis states all these points. The Spanish article also mentions the mine’s name and the company, which the Czech article doesn’t, but that’s not conflicting. The hypothesis doesn’t mention the company, so that’s okay. Are there any contradictions? The Spanish article mentions two people rescued, but the hypothesis says 27 died. Wait, the hypothesis says "leading to the death of 27 workers trapped inside the tunnel." The Spanish article says 27 died and two were rescued. The Czech article says 27 victims, which could include the two rescued? Wait, maybe not. If two were rescued, then the deaths would be 27, and the two rescued are alive. So the hypothesis says 27 workers died, which matches the Spanish article’s 27 deaths. The Czech article says "27 obětí" which is victims, which could be the 27 who died, but maybe the two rescued are not counted as victims? Or maybe the Czech article is using "obětí" as victims, which could include those who died and those rescued? Wait, in Czech, "obětí" is victims, usually meaning those who died. But the Spanish article says 27 died and two rescued. The Czech article says 27 victims, which might be the same 27, but the two rescued might not be considered victims. So the hypothesis says 27 workers died, which aligns with both articles. The Czech article says "27 obětí" which is victims, so that’s consistent. Also, the hypothesis mentions the fire caused by a short circuit, which both articles confirm. The mobilization of authorities and rescue teams is mentioned in both. The time is Saturday morning in both. The mine’s name and location are correct. Therefore, the hypothesis is fully supported by both articles. So the label should be entailed.

### v3

@@ E.2 LLM Answer Case
Qwen3-8B Reasoning Output Okay, let’s tackle this NLI task. So, the premise has two articles in Spanish and Czech, and the hypothesis is in English. I need to determine if the hypothesis is entailed, contradicted, or neutral based on the premises. First, I’$l$ read through the Spanish article. It mentions that on Saturday, there was a fire in a gold mine in Condesuyos province, caused by a short circuit 100 meters from the entrance. The mine is called Yanaquihua, known as Esperanza I. They say 27 people died, two were rescued. The regional government of Arequipa mentioned the short circuit caused the fire, risking workers’ lives. Local media said 27 trapped workers died of asphyxia. The Ministry of Interior confirmed the accident on Saturday, and police were there for rescue. The fire spread quickly through wooden structures, preventing escape. The mine belongs to Yanaquihua S.A.C., a company with small miners. Now the Czech article: Officials said the accident in the small gold mine in southern Peru resulted in 27 deaths. It happened on Saturday morning at the Yanaquihua mine in Condesuyos province, Arequipa department. A short circuit caused a fire inside the tunnel. The regional government confirmed 27 victims. The prosecutor Giovanni Matos said the information is correct, and police are there to find the dead, check safety for rescue teams. Regional government and interior ministry mobilized police, medics, and ambulances. It’s the highest number of deaths in a single mining accident since 2000. The company didn’t respond to a comment request. Hypothesis: The Yanaquihua gold mine in Condesuyos, Peru, had a fatal incident on Saturday morning due to a short circuit causing a fire, leading to 27 workers’ deaths trapped inside the tunnel, with local authorities and rescue teams mobilized. Check each part. The Spanish article says the fire was caused by a short circuit, leading to 27 deaths. The Czech article confirms the same: short circuit causing fire, 27 deaths. Both mention Saturday morning (the Spanish says "sábado" which is Saturday, and the Czech says "v sobotu ráno" which is Saturday morning). The mine is Yanaquihua in Condesuyos, Peru. The death count is 27, and both mention authorities and rescue teams being mobilized. The hypothesis states all these points. The Spanish article also mentions the mine’s name and the company, which the Czech article doesn’t, but that’s not conflicting. The hypothesis doesn’t mention the company, so that’s okay. Are there any contradictions? The Spanish article mentions two people rescued, but the hypothesis says 27 died. Wait, the hypothesis says "leading to the death of 27 workers trapped inside the tunnel." The Spanish article says 27 died and two were rescued. The Czech article says 27 victims, which could include the two rescued? Wait, maybe not. If two were rescued, then the deaths would be 27, and the two rescued are alive. So the hypothesis says 27 workers died, which matches the Spanish article’s 27 deaths. The Czech article says "27 obětí" which is victims, which could be the 27 who died, but maybe the two rescued are not counted as victims? Or maybe the Czech article is using "obětí" as victims, which could include those who died and those rescued? Wait, in Czech, "obětí" is victims, usually meaning those who died. But the Spanish article says 27 died and two rescued. The Czech article says 27 victims, which might be the same 27, but the two rescued might not be considered victims. So the hypothesis says 27 workers died, which aligns with both articles. The Czech article says "27 obětí" which is victims, so that’s consistent. Also, the hypothesis mentions the fire caused by a short circuit, which both articles confirm. The mobilization of authorities and rescue teams is mentioned in both. The time is Saturday morning in both. The mine’s name and location are correct. Therefore, the hypothesis is fully supported by both articles. So the label should be entailed.